OpenStudy (studentchaitrali):

1x1! +2x2!+3x3!+4x4!+....+nxn!

OpenStudy (anonymous):

What is this?

OpenStudy (freckles):

What do you want to do with this sum?

OpenStudy (studentchaitrali):

find the value

OpenStudy (studentchaitrali):

in terms of n

OpenStudy (anonymous):

Prove that 1*1! + 2*2! + 3*3! + .... + n*n! = (n + 1)! - 1. 3 steps in induction: a) Prove true for trivial case n = 1. b) Assume true for some n = k; this is the Inductive Hypothesis IH. c) Assuming IH, prove for n = k + 1. a) Prove for n = 1: 1*1! = 1. b) Assume 1*1! + 2*2! + 3*3! + .... + n*n!=(n + 1)! - 1 (IH). c) Now we want to show 1*1! + 2*2! + 3*3! + .... + n*n! + (n + 1)*(n + 1)! = (n + 2)!-1. LHS = [1*1! + 2*2! + 3*3! + .... + n*n!] + (n + 1)*(n + 1)! = (n + 1)! - 1 + (n + 1)*(n + 1)! (replace [ ] by IH) = (n + 1)![1 + n + 1] - 1 = (n + 1)!*(n + 2) - 1 = (n + 2)! - 1 = RHS, since the recursive rule for factorials is (m + 1)*m! = (m + 1)!.

OpenStudy (freckles):

consider using the following: \[n \cdot n! =(n+1-1) \cdot n!=(n+1) \cdot n!-1 \cdot n! \\ n \cdot n! =(n+1) n!-n!=(n+1)!-n! \\ n \cdot n!=(n+1)!-n!\] you would rewrite each term for example: \[3 \cdot 3!=4!-3!\]